A standing wave on a string has a frequency of 100 Hz, a wavelength of 0.04 m, and an amplitude of 2 mm. The transverse velocity of the point x = 0.02 m, is:


1
Expert's answer
2021-12-20T10:31:06-0500

The transverse velocity is


"v(x,t)=-Aω \cos (k x - ω t + φ),\\
v(x,t)=-A(2\pi f) \cos [(2\pi/\lambda) x - (2\pi f) t+0],\\
v(0.02,t)=1.26\cos(3.14-628t)."